calculate simultaneous solubility of AgCNS and AgBr in water Ksp=1e-12 and 5e-13??


 methord please

Grade:11

2 Answers

Abhishek sharma
6 years ago
AbhigahahhahahaHahaha shajw sjw s ajw amWhwv w s wks e gphhezss ssvus9hodds diohpe d jp3d 3d ggfjm nbvb jbfg m yb m vesacb h
Abhishek sharma
6 years ago
AgCNS----->Ag+ + CNS- (X) (X)AgBr---------->Ag+ + Br- (Y) (Y) Ksp(AgCNS)= 1×10^-12 = (X+Y)X ---(1)Ksp(AgBr)=5×10^-13 = (X+Y)Y ---(2) Solving (1) and (2)[AgCNS]=X= 8.16×10^-7[AgBr]=Y= 4.08×10^-7 Which is the required answer